# American Option Pricing Under Time-Varying Rough Volatility: A Signature-Based Hybrid Framework

### v1

@@ 1 Introduction
American options permit exercise at any time before expiration, so their value depends on their entire price path rather than a single terminal payoff ([8]). Pricing such contracts constitutes an optimal–stopping problem, whose complexity increases substantially when realistic market features, such as stochastic volatility or roughness, are introduced. The fair value of an American option is given by:
$V_{t}=\sup_{\tau\in\mathcal{T}_{t,T}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\tau-t)}\,\phi(S_{\tau})\,\middle|\,\mathcal{F}_{t}\right]$ (1)
@@
    $\phi(S_{\tau})=\begin{cases}(K-S_{\tau})^{+}&\text{for a put option}\\ (S_{\tau}-K)^{+}&\text{for a call option}\end{cases}$ (2)
@@
Precise valuation matters for hedging, capital planning, and regulatory compliance. Yet in practice, traders often resort to ad-hoc adjustments and heuristic early-exercise premia111 Traders often adjust European option prices with informal “early-exercise premia” to approximate American values. These are rule-of-thumb add-ons used when models are unreliable—especially during volatility spikes or illiquid periods. when classical models fail under market stress.
As shown in Black and Scholes (1973) and Merton (1973), the classical framework assumes constant volatility and memoryless price variation—assumptions that conflict with empirical evidence of volatility clustering, leverage effects, and long-range dependence [6, 18]. While lattice and finite-difference methods remain operable in these settings, [^2] Lattice and finite-difference methods numerically approximate option prices by discretizing time and state variables. These methods can incorporate early exercise features and certain types of path dependence, but often struggle to capture complex dynamics like stochastic volatility or rough paths. their outputs frequently diverge from observed market prices.
@@
A growing body of evidence shows that volatility behaves like fractional Brownian motion with a Hurst parameter below one half—so-called ’rough volatility,’ as described by Gatheral, Jaisson, and Rosenbaum (2018) and Bennedsen, Lunde, and Pakkanen (2016) [9, 5]. The Hurst parameter $H\in(0,1)$ measures the regularity of a path: values below one half correspond to anti-persistent behavior and rapidly mean-reverting fluctuations, while values above one half indicate smoother trends. In markets, empirical estimators such as the rescaled-range statistic, originally developed by Hurst (1951) and Mandelbrot and Wallis (1969) show that volatility roughness is typically below one half but drifts upward during stable periods, reflecting a smoother environment [11, 16]. Capturing this time-varying behavior is essential for American option pricing models to remain realistic and responsive to changing market regimes.
@@
A recent breakthrough by Bayer, Pelizzari, and Zhu [3] reframes the American option pricing problem under rough volatility using rough-path theory. Their method encodes volatility trajectories using truncated path signatures and applies kernel or linear regression to estimate continuation values and construct martingale controls, yielding tight duality gaps. Signatures offer a universal and theoretically complete representation of non-Markovian trajectories [14], but the approach presented by Bayer et al. remains limited in practice. First, the original model assumes constant parameters throughout, including a fixed Hurst parameter, despite empirical evidence that volatility roughness evolves over time. Second, it relies on deep neural networks and exact signature kernels, which scale at least quadratically with path length and require specialized hardware to execute efficiently.
@@
The present work proposes a relatively lightweight extension that removes these obstacles while preserving the theoretical foundation of signature representations. A gradient-boosted tree ensemble [7] is trained on sequential rolling windows of volatility signatures, short overlapping time slices of recent market behavior—to forecast the evolving Hurst path $H(t)$. This forecast informs a simple rule: when the average $H(t)$ is below 0.5, the contract is priced with the rough Bergomi engine; if it is at or above 0.5, we switch to a calibrated Heston engine, which runs more efficiently in smoother regimes and reduces the duality gaps observed under a purely rough specification.
@@
Beyond dynamic roughness adaptation, this work introduces further practical enhancements. Signature kernel evaluation is accelerated using Random Fourier Features (RFFs) [23], in contrast to Bayer et al. who rely on exact signature kernels that scale poorly and require specialized hardware. This approximates the kernel in linear time and enables our model to handle longer volatility paths ($N=252$) without significant loss in accuracy. In addition, key parameters are computed from daily at-the-money (ATM) options data instead of relying on fixed calibration, allowing for adaptability to real market data. To improve granularity, Bayer’s original codebase was modified to support pricing with calendar days rather than fractional years. These enhancements reflect not only a re-implementation but a redesign, intended to bring signature-based methods closer to real deployment.
@@
Overall, our framework extends Bayer et al. by adapting to time-varying roughness, providing a more accurate and flexible architecture for pricing American options under non-smooth volatility. However, due to computational constraints, extensive empirical validation remains challenging. Bayer et al. (2025) note that duality gaps are sensitive to truncation level $K$, path length $N$, and Monte Carlo size $M$. In this implementation, we use $K=3$, $N=252$, and $M=2^{15}$, which reflects hardware limitations rather than an optimized configuration.
@@ 2.1 Overview
A multi-step gradient-boosted tree ensemble then forecasts Hurst parameter values for each of the next $\tau$ trading days; the resulting path of predicted $H$ values drives an automatic switch between two stochastic-volatility engines. When the forecast average remains below one half, simulations are generated with the rough Bergomi model, whose fractional kernel is documented in high-frequency data by Bayer, Friz, and Gatheral (2016) [4] and is defined explicitly in Equation (7).If the forecast crosses the threshold (0.5), the framework reverts to a calibrated Heston specification that is known to capture smoother regimes efficiently, as shown by Heston (1993) [10]. Each simulated path is lifted into a truncated signature of degree three and level three; signatures form a universal, algebraic feature map for non-Markovian dynamics [15]. Lower and upper price bounds are then computed by the primal–dual machinery introduced by Bayer, Pelizzari and Zhu [3]. Because every feature of the Hurst parameter (estimation, forecasting, simulation, and regression) can be replaced independently, the pipeline accommodates future methodological advances without wholesale redesign.
@@ 2.3 Volatility Engines and Calibration
 [^3] These formulas allow for fast, daily re-estimation using only price and implied volatility data: (i) $\rho$ is estimated as the sample correlation between daily log returns and changes in implied volatility. This serves as a simple proxy for the leverage effect, though it may be noisy at daily frequency [1]. (ii) The formula for $\eta$ is derived by rearranging the variance scaling relationship described in Gatheral et al. (2018): $\operatorname{Var}[\Delta\log\sigma]\approx\eta^{2}(\Delta t)^{2H}$, which yields $\eta=\operatorname{std}(\Delta\log\sigma)/(\Delta t)^{H}$ under the assumption that log volatility increments follow a rough path with Hurst parameter $H$ [9]. We emphasize that $\eta$ is calibrated using the most recent historical Hurst estimate $H_{t}$ (as defined in Equation (4)), not the average forecast $\bar{H}_{t}$, in order to avoid lookahead bias. Calibration should rely only on currently available information to maintain forward-valid modeling. (iii) $\xi_{0}$ is set equal to the square of the current ATM implied volatility, following common practice in rough volatility models [17, 22].
Rough Bergomi Engine. When forecasted market roughness is high ($\bar{H}_{t}<0.5$), asset and variance paths are simulated under the rough Bergomi system [3]:
@@
We adapt the rough Bergomi specification of Bayer, Friz, and Gatheral (2016) [4], but extend it to incorporate a time-varying Hurst path. The kernel exponent in the stochastic integral is updated at each time increment using the forecast $\widehat{H}_{u}$, allowing the integrand to adapt to evolving roughness. However, the normalization term $\frac{1}{2}\eta^{2}s^{2\widehat{H}_{t}}$ assumes $\widehat{H}_{t}$ is held constant over $[0,s]$ for tractability. This hybrid treatment preserves local roughness dynamics in simulation while simplifying variance scaling.
@@
The smoother regime uses the classical Heston model [10], in which variance follows a mean-reverting square-root diffusion. The correlation $\rho$ introduces leverage effects, and the model calibrates efficiently in calm markets. Regime switching is implemented through two modular simulation classes: DynamicHurstrBergomi, which applies a time-varying Volterra kernel using a Hurst forecast path, and DynamicHeston, which uses the classical dynamics above. Both classes share a common interface, enabling seamless substitution within the pricing pipeline.
@@ 2.4 Efficient Signature Regression via Kernel Approximation
Lyons (2014) shows that path signatures separate trajectories up to tree‐like equivalence and therefore provide a universal basis for non-Markovian functionals [14]. For a time-augmented volatility path $Z:[0,T]\to\mathbb{R}^{d}$ we retain the truncated signature of degree three and level three, $\operatorname{Sig}_{0,T}(Z)$. Higher degrees increase expressive power, but empirical work on short-dated equity options finds little benefit beyond degree three once sampling variance is considered [3].
@@
  • $\|x-y\|^{2}$ is their squared Euclidean distance.
@@
To restore scalability, we adopt Random Fourier Features (RFF), which approximate any shift-invariant kernel using a finite trigonometric map [23]:
@@ 2.5 Understanding the Results: Primal–Dual Bounds with Four Signature Variants
Upper and lower price estimates follow the primal–dual machinery of Bayer, Pelizzari, and Zhu [3], which builds on least-squares Monte Carlo [13] and the martingale dual of Rogers [24]. On a daily exercise mesh $t_{0}<\dots<t_{N_{1}}=T$ the primal recursion sets
@@
  • Extended Linear Signature – linear terms plus selected quadratic cross-products, following Pérez-Arribas [21].
  • Deep Log-Signature – three-layer perceptron on log-signatures, after Bai et al. [2].
  • Deep Kernel (RFF) – Random Fourier map on raw signatures followed by kernel ridge regression [23].
Running all four in parallel shows clear trade-offs. The Deep Kernel variant narrows duality gaps the most while adding less than 20 % to runtime. The Deep Log-Signature can diverge, inflating upper bounds—an instability echoed in recent deep-signature studies. These results confirm that lightweight RFF maps capture much of the non-linear benefit of a full Gaussian signature kernel without its quadratic cost, consistent with other kernel-scalability work [12].
@@ 3.1 Conceptual and Practical Contributions
This paper extends signature‐based American option pricing by integrating a forward-looking forecast of volatility roughness. Instead of assuming a fixed Hurst parameter—as in Bayer, Pelizzari, and Zhu [3]—we estimate and forecast roughness dynamically with an ensemble of XGBoost regressors trained on the most recent realized values. Based on this forecast, the model selects a volatility engine (rough Bergomi or Heston) once per contract and uses it to simulate all paths over the remaining lifetime. Empirical studies show that the average Hurst parameter $\bar{H}_{t}$ tends to fall below $0.5$ during stress and drift toward Brownian levels in calm periods [9, 5]. Routing simulations on the basis of this forecast embeds documented regime shifts directly into the pricing engine.
@@ 3.2 Relationship to Prior Work and Intended Contributions
Fixed-roughness architectures. One stream assumes volatility roughness is fixed and known. Bayer, Pelizzari, and Zhu [3] exemplify this approach by pricing American puts on simulated rough Bergomi paths with constant Hurst parameter $H=0.07$. Their deep-signature networks estimate continuation values and construct martingale controls, achieving narrow duality gaps on synthetic data. However, this framework assumes a static roughness regime—an assumption contradicted by empirical findings.
Empirical evidence of evolving roughness. Studies by Gatheral, Jaisson, and Rosenbaum (2018) and Bennedsen, Lunde, and Pakkanen (2016) demonstrate that roughness varies over time: falling during market stress and drifting toward Brownian levels in calm periods [9, 5]. This suggests that a fixed-$H$ simulator may mismatch real-world dynamics when volatility regimes shift.
@@ 3.3 Limitations and Methodological Caveats
Despite this speed-up, the resulting duality gaps remain non-negligible—echoing findings by Bayer et al. [3], who note that tighter bounds often require many more paths, deeper signatures, or richer feature maps. These remain difficult to achieve on CPU-only hardware.
Variance from random features and deep log-signatures. RFFs introduce stochastic variance that grows with $D$; empirical tuning suggests $D=128$ balances bias and variance, but formal convergence guarantees under rough-path inputs are still open. Likewise, the deep log-signature variant can destabilize the dual optimiser, inflating upper bounds when learning-rate or weight initialisation interacts poorly with highly irregular paths. Robust regularisation and cross-validation remain essential future steps.
Summary. The architecture offers a pragmatic, modular path toward option pricing under time-varying roughness: it adapts to shifting regimes, slashes computational overhead relative to full kernel methods, and preserves path-wise expressiveness. Yet it also inherits noise from roughness estimation, relies on shallow calibration, and is sensitive to simulation budget and feature variance. Addressing these caveats—through better roughness estimators, richer calibration targets, larger simulation budgets, or more stable optimisation—defines the roadmap for future work.

### v2

@@ 1 Introduction
American options permit exercise at any time before expiration, so their value depends on their entire price path rather than a single terminal payoff (CME Group 2025). Pricing such contracts constitutes an optimal–stopping problem, whose complexity increases substantially when realistic market features, such as stochastic volatility or roughness, are introduced. The fair value of an American option is given by:
$V_{t}=\sup_{\tau\in\mathcal{T}_{t,T}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\tau-t)}\,\phi(S_{\tau})\,\middle|\,\mathcal{F}_{t}\right],$ (1)
where:
@@
    $\phi(S_{\tau})=\begin{cases}(K-S_{\tau})^{+}&\text{for a put option},\\ (S_{\tau}-K)^{+}&\text{for a call option},\end{cases}$ (2)
@@
Precise valuation matters for hedging, capital planning, and regulatory compliance. Yet in practice, traders often resort to ad-hoc adjustments and heuristic early-exercise premia [^1] Traders often adjust European option prices with informal “early-exercise premia” to approximate American values. These are rule-of-thumb add-ons used when models are unreliable—especially during volatility spikes or illiquid periods. when classical models fail under market stress.
As shown in Black and Scholes (1973) and Merton (1973), the classical framework assumes constant volatility and memoryless price variation—assumptions that conflict with empirical evidence of volatility clustering, leverage effects, and long-range dependence. While lattice and finite-difference methods remain operable in these settings, [^2] Lattice and finite-difference methods numerically approximate option prices by discretizing time and state variables. These methods can incorporate early exercise features and certain types of path dependence, but often struggle to capture complex dynamics like stochastic volatility or rough paths. their outputs frequently diverge from observed market prices.
@@
A growing body of evidence shows that volatility behaves like fractional Brownian motion with a Hurst parameter below one half—so-called “rough volatility,” as described by Gatheral et al. (2018) and Bennedsen et al. (2016). The Hurst parameter $H\in(0,1)$ measures the regularity of a path: values below one half correspond to anti-persistent behavior and rapidly mean-reverting fluctuations, while values above one half indicate smoother trends. In markets, empirical estimators such as the rescaled-range statistic, originally developed by Hurst (1951) and Mandelbrot and Wallis (1969), show that volatility roughness is typically below one half but drifts upward during stable periods, reflecting a smoother environment. Capturing this time-varying behavior is essential for American option pricing models to remain realistic and responsive to changing market regimes.
@@
A recent breakthrough by Bayer et al. (2025) re-frames the American option pricing problem under rough volatility using rough-path theory. Their method encodes volatility trajectories using truncated path signatures and applies kernel or linear regression to estimate continuation values and construct martingale controls, yielding tight duality gaps. Signatures offer a universal and theoretically complete representation of non-Markovian trajectories, but the approach presented by Bayer et al. (2025) remains limited in practice. First, the original model assumes constant parameters throughout, including a fixed Hurst parameter, despite empirical evidence that volatility roughness evolves over time. Second, it relies on deep neural networks and exact signature kernels, which scale at least quadratically with path length and require specialized hardware to execute efficiently.
@@
The present work proposes a relatively lightweight extension that removes these obstacles while preserving the theoretical foundation of signature representations. A gradient-boosted tree ensemble is trained on sequential rolling windows of volatility signatures—short overlapping time slices of recent market behavior—to forecast the evolving Hurst path $H(t)$. This forecast informs a simple rule: when the average $H(t)$ is below 0.5, the contract is priced with the rough Bergomi engine; if it is at or above 0.5, we switch to a calibrated Heston engine, which runs more efficiently in smoother regimes and reduces the duality gaps observed under a purely rough specification.
@@
Beyond dynamic roughness adaptation, this work introduces further practical enhancements. Signature-kernel evaluation is accelerated using Random Fourier Features (RFFs) (Rahimi and Recht 2007), in contrast to Bayer et al. (2025) who rely on exact signature kernels that scale poorly and require specialized hardware. This approximates the kernel in linear time and enables our model to handle longer volatility paths ($N=252$) without significant loss in accuracy. In addition, key parameters are computed from daily at-the-money (ATM) options data instead of relying on fixed calibration, allowing for adaptability to real market data. To improve granularity, Bayer’s original codebase was modified to support pricing with calendar days rather than fractional years. These enhancements reflect not only a re-implementation but a redesign, intended to bring signature-based methods closer to real deployment.
@@
Overall, our framework extends Bayer et al. (2025) by adapting to time-varying roughness, providing a more accurate and flexible architecture for pricing American options under non-smooth volatility. However, due to computational constraints, extensive empirical validation remains challenging. Bayer et al. (2025) note that duality gaps are sensitive to truncation level $K$, path length $N$, and Monte Carlo size $M$. In this implementation, we use $K=3$, $N=252$, and $M=2^{15}$, which reflects hardware limitations rather than an optimized configuration.
@@ 2.1 Overview
A multi-step gradient-boosted tree ensemble then forecasts Hurst parameter values for each of the next $\tau$ trading days; the resulting path of predicted $H$ values drives an automatic switch between two stochastic-volatility engines. When the forecast average remains below one half, simulations are generated with the rough Bergomi model, whose fractional kernel is documented in high-frequency data by Bayer, Friz, and Gatheral (2016) and is defined explicitly in Equation (7). If the forecast crosses the threshold (0.5), the framework reverts to a calibrated Heston specification that is known to capture smoother regimes efficiently, as shown by Heston (1993). Each simulated path is lifted into a truncated signature of degree three and level three; signatures form a universal, algebraic feature map for non-Markovian dynamics (Lyons et al. 2014). Lower and upper price bounds are then computed by the primal–dual machinery introduced by Bayer, Pelizzari and Zhu (2025). Because every feature of the Hurst parameter (estimation, forecasting, simulation, and regression) can be replaced independently, the pipeline accommodates future methodological advances without wholesale redesign.
@@ 2.3 Volatility Engines and Calibration
 [^3] These formulas allow for fast, daily re-estimation using only price and implied volatility data: (i) $\rho$ is estimated as the sample correlation between daily log returns and changes in implied volatility. This serves as a simple proxy for the leverage effect, though it may be noisy at daily frequency (Aït-Sahalia et al. 2013). (ii) The formula for $\eta$ is derived by rearranging the variance scaling relationship described in Gatheral et al. (2018): $\operatorname{Var}[\Delta\log\sigma]\approx\eta^{2}(\Delta t)^{2H}$, which yields $\eta=\operatorname{std}(\Delta\log\sigma)/(\Delta t)^{H}$ under the assumption that log volatility increments follow a rough path with Hurst parameter $H$ (Gatheral et al. 2018). We emphasize that $\eta$ is calibrated using the most recent historical Hurst estimate $H_{t}$ (as defined in Equation (4)), not the average forecast $\bar{H}_{t}$, in order to avoid lookahead bias. Calibration should rely only on currently available information to maintain forward-valid modeling. (iii) $\xi_{0}$ is set equal to the square of the current ATM implied volatility, following common practice in rough volatility models (McCrickerd 2017; Quantitative Finance Stack Exchange 2021).
Rough Bergomi Engine. When forecasted market roughness is high ($\bar{H}_{t}<0.5$), asset and variance paths are simulated under the rough Bergomi system (Bayer et al. 2025):
@@
We adapt the rough Bergomi specification of Bayer, Friz, and Gatheral (2016), but extend it to incorporate a time-varying Hurst path. The kernel exponent in the stochastic integral is updated at each time increment using the forecast $\widehat{H}_{u}$, allowing the integrand to adapt to evolving roughness. However, the normalization term $\frac{1}{2}\eta^{2}s^{2\widehat{H}_{t}}$ assumes $\widehat{H}_{t}$ is held constant over $[0,s]$ for tractability. This hybrid treatment preserves local roughness dynamics in simulation while simplifying variance scaling.
@@
The smoother regime uses the classical Heston model (Heston 1993), in which variance follows a mean-reverting square-root diffusion. The correlation $\rho$ introduces leverage effects, and the model calibrates efficiently in calm markets. Regime switching is implemented through two modular simulation classes: DynamicHurstrBergomi, which applies a time-varying Volterra kernel using a Hurst forecast path, and DynamicHeston, which uses the classical dynamics above. Both classes share a common interface, enabling seamless substitution within the pricing pipeline.
@@ 2.4 Efficient Signature Regression via Kernel Approximation
Lyons (2014) shows that path signatures separate trajectories up to tree-like equivalence and therefore provide a universal basis for non-Markovian functionals. For a time-augmented volatility path $Z:[0,T]\to\mathbb{R}^{d}$ we retain the truncated signature of degree three and level three, $\operatorname{Sig}_{0,T}(Z)$. Higher degrees increase expressive power, but empirical work on short-dated equity options finds little benefit beyond degree three once sampling variance is considered (Bayer et al. 2025).
@@
  • $\|x-y\|^{2}$ is their squared Euclidean distance,
@@
To restore scalability, we adopt Random Fourier Features (RFF), which approximate any shift-invariant kernel using a finite trigonometric map (Rahimi and Recht 2007):
@@ 2.5 Understanding the Results: Primal–Dual Bounds with Four Signature Variants
Upper and lower price estimates follow the primal–dual machinery of Bayer, Pelizzari, and Zhu (2025), which builds on least-squares Monte Carlo (Longstaff and Schwartz 2001) and the martingale dual of Rogers (Rogers 2002). On a daily exercise mesh $t_{0}<\dots<t_{N_{1}}=T$ the primal recursion sets
@@
  • Extended Linear Signature – linear terms plus selected quadratic cross-products, following Pérez-Arribas (Pérez Arribas 2018).
  • Deep Log-Signature – three-layer perceptron on log-signatures, after Bai et al. (Bai et al. 2023).
  • Deep Kernel (RFF) – Random Fourier map on raw signatures followed by kernel ridge regression (Rahimi and Recht 2007).
Running all four in parallel shows clear trade-offs. The Deep Kernel variant narrows duality gaps the most while adding less than 20% to runtime. The Deep Log-Signature can diverge, inflating upper bounds—an instability echoed in recent deep-signature studies. These results confirm that lightweight RFF maps capture much of the non-linear benefit of a full Gaussian signature kernel without its quadratic cost, consistent with other kernel-scalability work (Le et al. 2013).
@@ 3.1 Conceptual and Practical Contributions
This paper extends signature-based American option pricing by integrating a forward-looking forecast of volatility roughness. Instead of assuming a fixed Hurst parameter—as in Bayer, Pelizzari, and Zhu (2025)—we estimate and forecast roughness dynamically with an ensemble of XGBoost regressors trained on the most recent realized values. Based on this forecast, the model selects a volatility engine (rough Bergomi or Heston) once per contract and uses it to simulate all paths over the remaining lifetime. Empirical studies show that the average Hurst parameter $\bar{H}_{t}$ tends to fall below $0.5$ during stress and drift toward Brownian levels in calm periods. Routing simulations on the basis of this forecast embeds documented regime shifts directly into the pricing engine.
@@ 3.2 Relationship to Prior Work and Intended Contributions
Fixed-roughness architectures. One stream assumes volatility roughness is fixed and known. Bayer, Pelizzari, and Zhu (2025) exemplify this approach by pricing American puts on simulated rough Bergomi paths with constant Hurst parameter $H=0.07$. Their deep-signature networks estimate continuation values and construct martingale controls, achieving narrow duality gaps on synthetic data. However, this framework assumes a static roughness regime—an assumption contradicted by empirical findings.
Empirical evidence of evolving roughness. Studies by Gatheral, Jaisson, and Rosenbaum (2018) and Bennedsen, Lunde, and Pakkanen (2016) demonstrate that roughness varies over time: falling during market stress and drifting toward Brownian levels in calm periods. This suggests that a fixed-$H$ simulator may mismatch real-world dynamics when volatility regimes shift.
@@ 3.3 Limitations and Methodological Caveats
Despite this speed-up, the resulting duality gaps remain non-negligible—echoing findings by Bayer et al. (2025), who note that tighter bounds often require many more paths, deeper signatures, or richer feature maps. These remain difficult to achieve on CPU-only hardware.
Variance from random features and deep log-signatures. RFFs introduce stochastic variance that grows with $D$; empirical tuning suggests $D=128$ balances bias and variance, but formal convergence guarantees under rough-path inputs are still open. Likewise, the deep log-signature variant can destabilize the dual optimizer, inflating upper bounds when learning-rate or weight initialization interacts poorly with highly irregular paths. Robust regularization and cross-validation remain essential future steps.
Summary. The architecture offers a pragmatic, modular path toward option pricing under time-varying roughness: it adapts to shifting regimes, slashes computational overhead relative to full kernel methods, and preserves path-wise expressiveness. Yet it also inherits noise from roughness estimation, relies on shallow calibration, and is sensitive to simulation budget and feature variance. Addressing these caveats—through better roughness estimators, richer calibration targets, larger simulation budgets, or more stable optimization—defines the roadmap for future work.